# Theory of excitonic order in kagome metal ScV6Sn6

### v2

@@ Abstract
We argue that kagome metals can feature an excitonic condensate of unconventional nature. Studying the recently discovered variant ScV6Sn6, we identify electron and hole pockets due to a pair of van Hove singularities (vHS) close to the Fermi level, with an approximate spectral particle-hole symmetry at low energies. A significant fraction of the Fermi level density of states away from the vHS is removed by the onset of high temperature charge density wave order, and makes the bands more two dimensional, setting the stage for the formation of excitons. We develop a two-orbital minimal model which captures these features, along with the sublattice support of the wavefunctions, and find $s$-wave or $d$-wave excitons depending on the microscopic interaction parameters – the latter of which exhibits either a charge nematic or time-reversal symmetry breaking flux order depending on strain, offering an explanation of recent STM and transport experiments. The appearance of particle-type and hole-type vHS, and the excitonic resonance associated with it, may be a common thread to understanding nematicity and time-reversal symmetry breaking in kagome metals.
@@
Figure 1: Excitonic orbital order in 166 kagome metal ScV6Sn6. The crystal structure of ScV6Sn6: (a) standard and (b) top-view of the pristine unit cell (V = red, Sc = violet, Sn = blue). (c) The presence of oppositely dispersing vHS results in a weak coupling instability towards an excitonic state which hybridizes the two bands, producing either flux (d, left) or nematic order (d, right).
@@
In this Letter, we develop a theory of excitonic resonances in kagome metals. Exploiting the abundance of vHS around pristine filling in kagome metals, we consider the case where opposite concavity vHS occur at a similar distance from the Fermi level. A properly chosen charge order can effectively convert a kagome metal to a ‘kagome semimetal’ at low energies, by gapping out a large fraction of complementary Fermi level density of states while leaving the vHS largely unspoiled. We show that this holds for recently discovered variant ScV6Sn6 Arachchige2022 , in contrast to $A$V3Sb5 in which the CDW develops within the same orbitals as a possible excitonic state Scammell2023 . Developing a minimal model for this material, we find $s$-wave and $d$-wave excitonic states as a function of interaction parameters, the latter of which offers an explanation of the nematic state recently seen in STM Jiang2023hove . The excitonic scenario we develop for ScV6Sn6 provides a theoretical reconciliation of several experimental facts, and suggests novel signatures in optical spectroscopy.
@@
Figure 2: Ab-initio bandstructure in the pristine and CDW phase of ScV6Sn6. DFT bands above (red) and below (green) the CDW transition, with V and Sn orbital weights. For better visualization, the Sn weights are scaled by a factor four times greater than the V weights. The band crossing the Fermi level in the $k_{z}=1/2$ plane derives from the hybridization between $p_{z}$ orbitals of Sn atoms and V atoms; the associated spectral weight is gapped out by the CDW, producing a more two-dimensional electronic structure.
@@
The materials are significantly more three dimensional than the $A$V3Sb5 series, due to the hybridization with the out-of-plane Sn orbitals – as reflected by the three-dimensional character of the dispersion shown in Fig. 2. This is also borne out in the lower transport anisotropy $\rho_{c}/\rho_{ab}\approx 5$ Mozaffari2023 ; DeStefano2023 and relative difficulty of cleaving samples Kang2023 . Yet, the spectral weight associated to the $p$-orbitals of the Sn atoms is gapped out upon the onset of the CDW, resulting in a significantly more two-dimensional dispersion, as well as pushing the two vHS closer to the Fermi level, Fig. 2, also reflected in the Fermi surface plots shown in the Supplementary Material (SM). As the CDW primarily affects the Sn and Sc atoms, the V-based kagome bandstructure remains intact.
Hence, we can construct a minimal model of ScV6Sn6 inside the CDW phase consisting of two effective $d$ orbitals per site on a kagome bilayer, representing the twelve bands predominantly composed of the vanadium $d_{x^{2}-y^{2}}$/$d_{xy}$ and $d_{z^{2}}$ orbitals respectively, corresponding to the $A_{2g}$ and $A_{1g}$ representations of the site symmetry group respectively. Since the dispersion arising from vanadium orbitals varies minimally along the $k_{z}$ direction inside the CDW phase, we work with a model that neglects the hoppings along the $z$ direction. Moving from layer basis to the basis of mirror eigenstates $(|\text{top}\rangle\pm|\text{bottom}\rangle)/\sqrt{2}$, the bands decouple into four sets of kagome bands with distinct mirror eigenvalues; projecting onto the two which cross the Fermi level, we obtain
@@
where $\alpha=z,x$ indexes the two orbitals, $\sigma$ sublattice, $t_{x}=0.69$ eV is the in-plane hopping between the $d_{xy}$-like orbitals, $t_{z}=-0.34$ eV is the in-plane hopping between the $d_{z^{2}}$ orbitals, and additionally, the two sets of orbitals possess different onsite energies, $\varepsilon_{z}=-0.77$ eV and $\varepsilon_{x}=1.35$ eV. Comparison between DFT and tight-binding in Fig. 3 indicates good agreement  [^2] Ref. Jiang2023 has also discussed tight-binding minimal models for 166 and the related 1:1 compounds. The models considered there apply to the pristine phase including the Sn $p$ orbitals, whereas our present model neglects the role of the $p$ orbitals, which we justify with resort to the effect of the CDW on the Sn spectral weight (Fig. 2)..
A key purpose of our minimal model is to accurately characterize the two vHS located close to the Fermi level, producing electron-like and hole-like Fermi surfaces with an approximate emergent particle-hole symmetry, consistent with prior ab-initio studies Tuniz2023 and ARPES arpes ; kundu2024low which observe two oppositely-dispersing vHS. The kagome tight-binding model exhibits a novel wavefunction structure near the vHS – referred to as the sublattice interference effect. Namely, the wavefunctions possess support either on one sublattice (pure, or $p$-type, sublattice support) or two (mixed, or $m$-type, sublattice support). We note importantly that the two vHS near the Fermi level are both $p$-type – as illustrated in Fig. 4. The plot of the charge density obtained from DFT shows that near a given $M$-point, the wavefunction of both bands is concentrated entirely on one sublattice. An accurate reflection of the sublattice support plays an important role in our analysis of interaction effects.
@@
On point ($i$): the two vHS cross each other, forming a ‘kagome semimetal’; if this band-crossing lies directly at the Fermi level, then for arbitrarily small dimensionless Coulomb interaction $g$, there exists a non-zero critical temperature at which excitons condense  [^3] The origin of weak-coupling instability towards exciton condensation can be understood as ‘zero wavevector nesting’: as is well known, when the Fermi surface possesses the nesting property $\varepsilon(\bm{k})=-\varepsilon(\bm{k}+\bm{Q})$ for some wavevector $\bm{Q}$, the Fermi surface becomes unstable towards the formation of a spin or charge density wave $\langle c^{\dagger}_{\bm{k}+\bm{Q}}c_{\bm{k}}\rangle\neq 0$. In the presence of two Fermi surfaces $c$ and $d$ with $\varepsilon_{c}(\bm{k})=-\varepsilon_{d}(\bm{k})$, one finds the nesting condition satisfied with $\bm{Q}=0$, indicating an instability towards a state which hybridizes the two Fermi surfaces $\langle c^{\dagger}_{\bm{k}}d_{\bm{k}}\rangle\neq 0$.. If the vHS cross away from Fermi level, as in ScV6Sn6, then there is an energy cost $\delta\varepsilon$ to create particle-hole pairs, which then appear as gapped states with some lifetime. Upon including interactions, the exciton has a total energy $\delta\varepsilon-\varepsilon_{b}$, where $\varepsilon_{b}$ is the binding energy. For $\varepsilon_{b}>\delta\varepsilon$, condensation occurs; this requires a minimum coupling strength $g^{*}$ such that $\delta\varepsilon=\Lambda e^{-1/\sqrt{g^{*}}}$, where $\Lambda$ is an energy cutoff  [^4] The $e^{-1/\sqrt{g}}$ dependence, rather than the $e^{-1/g}$ typical of BCS-like instabilities, appears due to the enhanced density of states near a vHS.. For ScV6Sn6, the relatively small $\delta\varepsilon\approx 20$ meV arpes suggests a tendency towards condensation.
@@
with $\Pi_{\bm{k}}$ the particle-hole susceptibility (see the SM). We shift the vHS to lie exactly at the Fermi level and solve numerically for $\Phi_{\bm{k}}$ in the $s$- and $d$-wave like irreps, $\Gamma_{\Phi}=\{A_{1g},E_{2g}\}$. Note that due to the Perron-Frobenius theorem, the lowest energy bound state of a purely attractive interaction potential will always be the $A_{1g}$ state – a theorem well-noted in the context of superconducting instabilities scheurer2024mechanism . Similarly, a purely attractive potential, such as that arising from $U^{\prime}$, will always favor $A_{1g}$ excitons. The orbital exchange term $-2V^{\prime}<0$ generates repulsion for excitons on different sublattices – and hence at different vHS, due to the $p$-type sublattice support – facilitating the appearance of $E_{2g}$ order.
@@
Discussion. In this work we have developed a minimal model for ScV6Sn6 in the CDW phase, and proposed that the system hosts excitonic order, discussing some candidate states based on a minimal model and offering an explanation of recent experiments. We find both a fully gapped phase which preserves time-reversal symmetry, and a nematic state which breaks rotational symmetry and selectively gaps the vHS – both states being possibly relevant to experiment. STM studies on kagome surface terminations observe nematic order around $70$ K – clearly distinguished from the CDW, which is absent on the kagome surface Jiang2023hove . Quasiparticle interference imaging observes a gap at the vHS at two of the three sets of $M$-points at this temperature; the appearance of a gap at the crossing between two oppositely-dispersing bands is a direct signature of an excitonic state. The excitonic state may also be related to the transport signatures of a second transition within the CDW phase Mozaffari2023 , and weak spectral features seen near the vanadium vHS Chen2023 ; Yang2024unveiling .
@@
Our related theoretical work in Ref. Scammell2023 predicted the appearance of such states in kagome metal CsV3Sb5, where they are expected to appear as bright resonances inside the CDW gap, due to coexistence of excitons with CDW of the same orbital content. In the case of ScV6Sn6, excitons form in a parent bandstructure which is semimetallic, which makes observation of these resonances more challenging as they are expected to be broad due to coupling with the electronic continuum. Excitons in the former situation are similar to those in transition metal dichalcogenides wang2018excitons , whereas those in the latter are analogous to those in candidate excitonic insulator Ta2NiSe5, a zero-gap semiconductor lu2017zero .
In both the case of ScV6Sn6 and AV3Sb5, the coexisting metallic bands which do not form excitons comprise different orbital content, suggesting that matrix element effects may suppress the exciton scattering rate and permit a feasible resolution; we leave a theoretical analysis of this possibility to future work, but suggest experimental investigation is the best approach to answer these questions.
Acknowledgements
JI thanks Riccardo Comin, Zurab Guguchia, Xiong Huang, Yu-Xiao Jiang, Chenhao Jin, Asish Kundu, William Meier, and Abhay Pasupathy for helpful discussions and comments. AC and RT acknowledge useful conversations with Matteo Dürrnagel, Werner Hanke, Steve Kivelson, and Riccardo Sorbello. A.C. acknowledges support from PNRR MUR project PE0000023-NQSTI. A.C., D.D.S. and R.T. acknowledge the Gauss Centre for Supercomputing e.V. (https://www.gauss-centre.eu) for funding this project by providing computing time on the GCS Supercomputer SuperMUC-NG at Leibniz Supercomputing Centre (https://www.lrz.de).

### v3

# Theory of excitonic order in kagome metals ScV6Sn6 and LuNb6Sn6

@@ Abstract
We argue that kagome metals can feature an excitonic condensate of unconventional nature. Studying the recently discovered variants ScV6Sn6 and LuNb6Sn6 we identify electron and hole pockets due to a pair of van Hove singularities (vHS) close to the Fermi level, with an approximate spectral particle-hole symmetry. A significant fraction of the Fermi level density of states away from the vHS is removed by the onset of high temperature charge density wave order, and makes the bands more two-dimensional, setting the stage for the formation of excitons. We develop a two-orbital minimal tight-binding model of these materials which captures these features along with the sublattice support of the wavefunctions, and find $s$- or $d$-wave excitons depending on interaction parameters – the latter of which exhibits either charge nematicity or time-reversal symmetry breaking (TRSB) depending on strain, offering an explanation of recent STM and transport experiments. The presence of particle- and hole-type vHS, and the associated excitonic resonance, may be a common thread to understanding nematicity and TRSB in kagome metals.
@@
Figure 1: Excitonic orbital order in 166 kagome metals. The crystal structure of ScV6Sn6: (a) standard and (b) top-view of the pristine unit cell (V = red, Sc = violet, Sn = blue). (c) The presence of oppositely dispersing vHS results in a weak coupling instability towards an excitonic state which hybridizes the two bands, producing either nematic order (d, right) in which the density of states (red) breaks rotational symmetry, or flux order (d, left) with rotationally symmetric density of states but with spontaneous loop currents (arrows).
@@
In this Letter, we develop a theory of excitonic resonances in kagome metals. Exploiting the abundance of vHS around pristine filling, we consider the case where opposite concavity vHS occur at a similar distance from the Fermi level. A properly chosen charge order can effectively convert a kagome metal to a ‘kagome semimetal’ at low energies, by gapping out a large fraction of complementary Fermi level density of states while leaving the vHS largely unspoiled. We show that this holds for recently discovered variants ScV6Sn6 and LuNb6Sn6 Arachchige2022 ; ortiz2025stability , in contrast to $A$V3Sb5 in which the CDW develops within the same orbitals as a possible excitonic state Scammell2023 ; ingham2025vestigial . Developing a minimal model, we find $s$-wave and $d$-wave excitonic states as a function of interaction parameters, the latter of which offers an explanation of the nematic state recently seen in STM Jiang2023hove . The excitonic scenario we develop for ScV6Sn6 provides a theoretical reconciliation of several experimental facts, and suggests novel signatures in optical spectroscopy.
@@
Figure 2: Ab initio bandstructure in the pristine and CDW phase of ScV6Sn6. DFT bands above (red) and below (green) the CDW transition, with V and Sn orbital weights. For better visualization, the Sn weights are scaled by a factor four times greater than the V weights. The band crossing the Fermi level in the $k_{z}=1/2$ plane derives from the hybridization between $p_{z}$ orbitals of Sn atoms and V atoms; the associated spectral weight is gapped out by the CDW, producing a more two-dimensional electronic structure.
@@
Recently, a sister compound LuNb6Sn6 has been discovered ortiz2025stability which exhibits very similar phenomenology, including the competing CDW orders wang2025formation , orbital content and bandstructure. Many of our predictions apply similarly to this compound, and so we shall focus our attention on ScV6Sn6, presenting relevant details on LuNb6Sn6 in the Supplementary Material (SM).
These materials are significantly more three dimensional than the $A$V3Sb5 series, due to the hybridization with the out-of-plane Sn orbitals – as reflected by the three-dimensional character of the dispersion shown in Fig. 2. This is also borne out in the lower transport anisotropy $\rho_{c}/\rho_{ab}\approx 5$ Mozaffari2023 ; DeStefano2023 and relative difficulty of cleaving samples Kang2023 . Yet, the spectral weight associated to the $p$-orbitals of the Sn atoms is gapped out upon the onset of the CDW, resulting in a significantly more two-dimensional dispersion, as well as pushing the two vHS closer to the Fermi level, Fig. 2, also reflected in the Fermi surface plots shown in the SM. As the CDW primarily affects the Sn and Sc/Lu atoms, the V-based kagome bandstructure remains intact.
Hence, we can construct a minimal model of ScV6Sn6 and LuNb6Sn6 inside the CDW phase consisting of two effective $d$ orbitals per site on a kagome bilayer, representing the twelve bands predominantly composed of the vanadium $d_{x^{2}-y^{2}}$/$d_{xy}$ and $d_{z^{2}}$ orbitals respectively, corresponding to the $A_{2g}$ and $A_{1g}$ representations of the site symmetry group respectively. Since the dispersion arising from vanadium orbitals varies minimally along the $k_{z}$ direction inside the CDW phase, we work with a model that neglects the hoppings along the $z$ direction. Moving from layer basis to the basis of mirror eigenstates $(|\text{top}\rangle\pm|\text{bottom}\rangle)/\sqrt{2}$, the bands decouple into four sets of kagome bands with distinct mirror eigenvalues; projecting onto the two which cross the Fermi level, we obtain
@@
where $\alpha=z,x$ indexes the two orbitals, $\sigma$ sublattice, $t_{x}=0.69$ eV is the in-plane hopping between the $d_{xy}$-like orbitals, $t_{z}=-0.34$ eV is the in-plane hopping between the $d_{z^{2}}$ orbitals, and additionally, the two sets of orbitals possess different onsite energies, $\varepsilon_{z}=-0.77$ eV and $\varepsilon_{x}=1.35$ eV. Comparison between DFT and tight-binding in Fig. 3 indicates good agreement  [^2] Ref. Jiang2023 has also discussed tight-binding minimal models for 166 and the related 1:1 compounds. The models considered there apply to the pristine phase including the Sn $p$ orbitals, whereas our present model neglects the role of the $p$ orbitals, which we justify with resort to the effect of the CDW on the Sn spectral weight (Fig. 2).. Our model neglects the bands which cross the Fermi level near $K$ (Fig. 2), but these bands are poorly nested with the bands at $M$, and so do not contribute to the Fermi surface instabilities we describe. DFT and tight-binding fits for the two-orbital model of (Lu,Ho,Tb)Nb6Sn6 are given in the SM.
A key purpose of our minimal model is to accurately characterize the two vHS located close to the Fermi level, producing electron-like and hole-like Fermi surfaces with an approximate emergent particle-hole symmetry, consistent with prior ab initio studies Tuniz2023 and ARPES arpes ; kundu2024low which observe two oppositely-dispersing vHS. The kagome tight-binding model exhibits a novel wavefunction structure near the vHS – referred to as the sublattice interference effect. Namely, the wavefunctions possess support either on one sublattice (pure, or $p$-type, sublattice support) or two (mixed, or $m$-type, sublattice support). We note importantly that the two vHS near the Fermi level are both $p$-type – as illustrated in Fig. 4. The plot of the charge density obtained from DFT shows that near a given $M$-point, the wavefunction of both bands is concentrated entirely on one sublattice. An accurate reflection of the sublattice support plays an important role in our analysis of interaction effects.
@@
On point ($i$): the two vHS cross each other, forming a ‘kagome semimetal’; if this band-crossing lies directly at the Fermi level, then for arbitrarily small dimensionless Coulomb interaction $g$, there exists a non-zero critical temperature at which excitons condense  [^3] The origin of weak-coupling instability towards exciton condensation can be understood as ‘zero wavevector nesting’: as is well known, when the Fermi surface possesses the nesting property $\varepsilon(\bm{k})=-\varepsilon(\bm{k}+\bm{Q})$ for some wavevector $\bm{Q}$, the Fermi surface becomes unstable towards the formation of a spin or charge density wave $\langle c^{\dagger}_{\bm{k}+\bm{Q}}c_{\bm{k}}\rangle\neq 0$. In the presence of two Fermi surfaces $c$ and $d$ with $\varepsilon_{c}(\bm{k})=-\varepsilon_{d}(\bm{k})$, one finds the nesting condition satisfied with $\bm{Q}=0$, indicating an instability towards a state which hybridizes the two Fermi surfaces $\langle c^{\dagger}_{\bm{k}}d_{\bm{k}}\rangle\neq 0$.. If the vHS cross away from Fermi level, as in ScV6Sn6 and LuNb6Sn6, then there is an energy cost $\delta\varepsilon$ to create particle-hole pairs, which then appear as gapped states with some lifetime. Upon including interactions, the exciton has a total energy $\delta\varepsilon-\varepsilon_{b}$, where $\varepsilon_{b}$ is the binding energy. For $\varepsilon_{b}>\delta\varepsilon$, condensation occurs; this requires a minimum coupling strength $g^{*}$ such that $\delta\varepsilon=\Lambda e^{-1/\sqrt{g^{*}}}$, where $\Lambda$ is an energy cutoff  [^4] The $e^{-1/\sqrt{g}}$ dependence, rather than the $e^{-1/g}$ typical of BCS-like instabilities, appears due to the enhanced density of states near a vHS.. For ScV6Sn6, the relatively small $\delta\varepsilon\approx 20$ meV arpes suggests a tendency towards condensation  [^5] Note that DFT predicts $\delta\varepsilon\sim$ 50 meV, c.f. Fig. 4, overestimating the value of $\delta\varepsilon$ seen in experiment..
@@
with $\Pi_{\bm{k}}$ the particle-hole susceptibility (see the SM). Reaslistic values of $\delta\varepsilon$ have negligible effect on the phase diagram, since $\delta\varepsilon\ll U^{\prime},V^{\prime}$  [^6] We leave first-principles estimates of $U^{\prime}$ and $V^{\prime}$ to future work, but note that cRPA studies of 135 kagome metals with similar orbital content find $U^{\prime},V^{\prime}\sim$1-2 eV Wu2021 , and so we produce a phase diagram for a broad window of coupling constants with this order of magnitude., and so we shift the vHS to lie exactly at the Fermi level and solve numerically for $\Phi_{\bm{k}}$ in the $s$- and $d$-wave like irreps, $\Gamma_{\Phi}=\{A_{1g},E_{2g}\}$. Note that due to the Perron-Frobenius theorem, the lowest energy bound state of a purely attractive interaction potential will always be the $A_{1g}$ state – a theorem well-noted in the context of superconducting instabilities scheurer2024mechanism . Similarly, a purely attractive potential, such as that arising from $U^{\prime}$, will always favor $A_{1g}$ excitons. The orbital exchange term $-2V^{\prime}<0$ generates repulsion for excitons on different sublattices – and hence at different vHS, due to the $p$-type sublattice support – facilitating the appearance of $E_{2g}$ order. Undoing the particle-hole transformation, one finds that $V^{\prime}$ acts as an effective attractive interaction due to the unusual sublattice structure of the kagome lattice. Our results therefore reflect a subtle and new mechanism for nematicity in kagome systems, derived from a competition between onsite and neighbouring Coulomb interactions.
Note that our use of the Perron-Frobenius theorem applies to other particle-hole instabilities, including ferromagnetism; for this reason, most ordinary ferromagnets are $s$-wave, possessing trivial spatial structure. In recent years, great attention has been devoted to altermagnets, a kind of non $s$-wave ferromagnet smejkal2022beyond ; smejkal2022emerging . Our arguments imply that a purely repulsive Coulomb interaction cannot induce an altermagnetic phase via Fermi surface instability – some effective attraction is necessary, either from phonons or a non trivial sublattice structure, placing useful constraints on models of altermagnetism.
@@
Discussion. In this work we have developed a minimal model for 166 kagome metals in the $\sqrt{3}\times\sqrt{3}$ CDW phase, and proposed that this system hosts excitonic order, discussing a new mechanism for non $s$-wave orders and offering an explanation of recent experiments. We find both a fully gapped phase which preserves time-reversal symmetry, and a nematic state which breaks rotational symmetry and selectively gaps the vHS – both states being possibly relevant to experiment. Interestingly, our prediction of nematicity and possible flux order relied solely on the nature of the bandstructure and orbital content, which we have shown to apply also to LuNb6Sn6; we therefore predict the observation of nematicity, and unusual transport signatures in this sister compound.
STM studies on kagome surface terminations of ScV6Sn6 observe nematic order around $70$ K – clearly distinguished from the CDW, which is absent on the kagome surface Jiang2023hove . Quasiparticle interference imaging observes a gap at the vHS at two of the three sets of $M$-points at this temperature; the appearance of a gap at the crossing between two oppositely-dispersing bands is a direct signature of an excitonic state, distinct from signatures produced by other nematic states nab2024pomeranchuk . The fact that the nematic state affects the bandstructure at $M$ most strongly indicates that the crucial bands are those in our two orbital model, which neglects the bands near $K$. The excitonic state may also be related to the transport signatures of a second transition within the CDW phase Mozaffari2023 – near which nematic features have recently been seen farhang2025discovery – and weak spectral features seen near the vanadium vHS Chen2023 ; Yang2024unveiling .
@@
Our related theoretical work in Ref. Scammell2023 predicted the appearance of such states in kagome metal CsV3Sb5, where they are expected to appear as bright resonances inside the CDW gap, due to coexistence of excitons with CDW of the same orbital content. In the case of ScV6Sn6 and LuNb6Sn6, excitons form in a parent bandstructure which is semimetallic, which makes observation of these resonances more challenging as they are expected to be broad due to coupling with the electronic continuum. Excitons in the former situation are similar to those in transition metal dichalcogenides wang2018excitons , whereas those in the latter are analogous to those in candidate excitonic insulator Ta2NiSe5, a zero-gap semiconductor lu2017zero .
In both the cases of $A$V3Sb5 and ScV6Sn6/LuNb6Sn6, the coexisting metallic bands which do not form excitons comprise different orbital content, suggesting that matrix element effects may suppress the exciton scattering rate and permit a feasible resolution; we leave a theoretical analysis of this possibility to future work, but suggest experimental investigation is the best approach to answer these questions.
Acknowledgments
J.I. thanks Riccardo Comin, Zurab Guguchia, Xiong Huang, Yu-Xiao Jiang, Chenhao Jin, Asish Kundu, William Meier, and Abhay Pasupathy for helpful discussions and comments. J.I. is supported by NSF Career Award No. DMR-2340394. A.C. and R.T. acknowledge useful conversations with Matteo Dürrnagel, Werner Hanke, Steve Kivelson, and Riccardo Sorbello. A.C. acknowledges support from PNRR MUR project PE0000023-NQSTI. A.C., D.D.S. and R.T. acknowledge the Gauss Centre for Supercomputing e.V. (https://www.gauss-centre.eu) for funding this project by providing computing time on the GCS Supercomputer SuperMUC-NG at Leibniz Supercomputing Centre (https://www.lrz.de).